# Phase Selection and Analysis for Multi-frequency Multi-user RIS Systems Employing Subsurfaces in Correlated Ricean and Rayleigh Environments

### v1

@@ IV-C Special Case 2: Mean SNR for LoS $\mathbf{H}_{\mathrm{rb}}$ with Correlated Rayleigh UE Channels
$\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{f}_{k}]=\frac{N_{k}\pi||\mathbf{R}^{(k)1/2}_{\mathrm{d}}\mathbf{a}^{(k)}_{\mathrm{b}}||\sqrt{\beta^{(k)}_{\mathrm{d}}\beta^{(k)}_{\mathrm{rb}}\beta^{(k)}_{\mathrm{ur}}}}{4},$ (47)
@@
IV-D Insights from Analysis
From the analysis, the following insights are drawn. Equations (20) - (26), (30) (33) and (36) demonstrate the obvious properties that SNR improves with an increase in the number of RIS elements per user, BS elements and/or channel gain.
Both (33) and (36) include a Gaussian hypergeometric function, which increases monotonically with the final argument. The final arguments of both terms increase as the RIS spatial channel correlation increases, which occurs when RIS elements are placed closer together. In contrast, higher element spacing at the BS increases rate, due to the $||\mathbf{R}_{\mathrm{d}}\mathbf{a}_{\mathrm{b}}||$ component of (23). If elements are spaced too close together at the BS, this term decreases, as is shown in [13]. However, as the terms in (33) and (36) are quadratic, and thus typically the largest, a change in spatial correlation at the RIS is more influential to the overall SNR than a change in spatial correlation at the BS.
Therefore, the mean SNR is improved by typical factors such as large system dimensions and channel gain, and by the less obvious channel features of an uncorrelated $\mathbf{h}_{\mathrm{d}}$ channel and a highly correlated $\mathbf{h}_{\mathrm{ru}}$ channel.
@@ V Numerical Results
It is assumed that the RIS is fixed at 40 m away from the BS, at an angle of $\frac{\pi}{4}$ rad. Users are then dropped in a straight corridor of width $2w$ between BS and RIS, at a maximum distance of $r$ from the RIS with an exclusion zone of 1 m around the RIS. This region ensures that the path from the UE through the RIS significantly contributes to the total channel, and the effects of the RIS can be clearly seen. User clusters are considered to investigate the effect of people naturally congregating. Three types of user grouping were used in simulations. As seen in Fig. 2, layout A places all $K$ users in one cluster, where they are all in the same area spaced 1 m apart and layout B creates 2 clusters of $\frac{K}{2}$ users, with users spaced 1 m apart. Cluster locations are drawn from a uniform distribution. Layout C places each $K$ user randomly within the corridor, again using a uniform distribution. To ensure that results are general and not tied to a specific UE location, $10^{4}$ random user drops are generated per layout, each with $10^{6}$ replicates.
@@
The steering vectors $\mathbf{a}_{\mathrm{d}}$, $\mathbf{a}_{\mathrm{b}}$, $\mathbf{a}_{\mathrm{r}}$ and $\mathbf{a}_{\mathrm{ur}}$ correspond to the vertical uniform rectangular array (VURA) model and are given in [22]. Elements are arranged in a VURA at intervals of $d_{b}$ and $d_{r}$ wavelengths at the BS and RIS, respectively. $M_{x}$ and $N_{x}$ are the number of BS and RIS elements per row and $M_{z}$ and $N_{z}$ are the number of BS and RIS elements per column, such that $M=M_{x}M_{z}$ and $N=N_{x}N_{z}$. $\theta_{\mathrm{D}}$ and $\phi_{\mathrm{D}}$ are the elevation and azimuth angles of departure (AoDs) at the RIS and $\theta_{\mathrm{A}}$ and $\phi_{\mathrm{A}}$ are the corresponding elevation and azimuth angles of arrival (AoAs) at the BS. We assume the RIS is on a $\frac{\pi}{4}$ rad angle with respect to the BS, so $\phi_{\mathrm{D}}=\frac{5\pi}{4}$ rad and $\phi_{\mathrm{A}}=\frac{\pi}{4}$ rad. We also assume both are at the same height, so $\theta_{\mathrm{D}}=\theta_{\mathrm{A}}=\frac{\pi}{2}$ rad.
Without loss of generality we assume $\sigma^{2}=1$ and $E_{s}$ is selected in Sec V-A so that the mean output SNR for one UE is 5 dB in a system using the SD where $N=128$, $M=16$, $d_{r}=0.1$, $d_{b}=0.5$ and $\kappa_{d}=\kappa_{rb}=\kappa_{ur}=1$. In Sec V-B, we select $E_{s}$ so that the mean output SNR for one UE is 5 dB in a system using the SD where $N=128$, $M=16$, $d_{r}=0.1$, $d_{b}=0.5$, $\kappa_{\mathrm{ur}}=\kappa_{\mathrm{d}}=1$ and the RIS-BS channel is purely LoS, an important special case. These parameter values and definitions do not change throughout the results, unless specified otherwise.
@@
Fig 3 verifies the analytical results in Sec IV-A by showing the equivalence of $\mathbb{E}[\mathrm{SNR}]$ and the simulated mean SNR for the SD, and then compares this with the simulated mean SNR for the ISD. Two versions of the ISD were implemented. The first iterates through all subsurfaces once, as shown in Sec III-C. The second iterates through all subsurfaces and calculates the SNR of the system. It then repeats this process multiple times until the SNR of the system has converged, i.e. it increases by less than a specified tolerance from one iteration to the next. For these simulations, the tolerance was specified to be $10^{-4}$. We will refer to this method as the converged ISD (CISD).
The array parameters for Fig. 3 are $M=16$, $K=4$, $M_{x}=4$, $d_{b}=d_{r}=0.5$ and all K-factors are 1. Larger spacings between the RIS elements ($d_{r}=0.5$ rather than $d_{r}=0.1$ ) were used here to highlight the performance differences of the methods. Higher correlation leads to greater similarity in the SNR per user of the SD and CISD, as discussed below.
@@
Figure 3: Comparison of mean SNR for a randomly selected user for different subsurface phase selection methods while varying the number of RIS elements.
@@
It can be seen that the simulated mean SNR for the method outlined in Sec III-A is virtually indistinguishable from the $\mathbb{E}[\mathrm{SNR}]$ derived in Sec. IV-A, verifying that result. As expected, the mean SNR for all 3 methods increases with $N$, however the rate of improvement decreases as $N$ increases. This is expected, as it is well known that the RIS SNR can grow quadratically with $N$ [13] so that the dB relationship is logarithmic. The quadratic behaviour can be seen in (33), (43), and (48) where the $\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}]$ term contains the sum of $N_{k}(N_{k}-1)$ positive terms, giving a contribution of $\mathcal{O}(N_{k}^{2})$. The ISD and converged ISD both provide more benefit at higher $N$. At high $N$, more improvement can be made by iterating, as there are more elements to vary and hence more opportunities to improve the performance of the system. This improvement is more pronounced when the correlation between RIS elements is lower, as discussed below. The mean SNR for a range of $N$ at both RIS element spacings of $d_{r}=0.1$ and $0.5$ is shown below in Table I. The ISD performs better when correlation is lower. This may be due to the fact that there is more variability in the previously set subsurfaces creating the fixed channels for the iterative design. Hence, less correlation causes increased variation in the fixed channels which the iterative design can make use of.
TABLE I: Mean SNR for SD and CISD for different RIS element correlations
N Mean SNR (dB)
$\mathbf{d_{r}=0.1}$ $\mathbf{d_{r}=0.5}$
SD CISD SD CISD
$N=4$ $-3.24$ $-3.18$ $-3.19$ $-3.17$
$N=8$ $-1.05$ $-0.76$ $-2.52$ $-2.45$
$N=16$ $0.85$ $1.23$ $-1.08$ $-0.89$
$N=32$ $2.82$ $3.23$ $0.13$ $0.72$
$N=64$ $5.23$ $5.55$ $1.66$ $2.51$
$N=128$ $8.65$ $9.48$ $5.00$ $6.24$
@@
Multiplies $KM\!+\!N$ $NK\!+\!n\!\Big{(}\!\frac{(K\!-\!1)!MN}{K}$$\times\!\!\left(\!\frac{N}{K}\!+\!1\!\right)\!+\!KM\!+\!N\!\Big{)}$ $NK\!+\!n\!\Big{(}\!\frac{(K\!-\!1)!MN}{K}$$\times\!\!\left(\!\frac{N}{K}\!+\!1\!\right)\!+\!KM\!(2\!+\!N(N\!+\!1)\!)\!+\!N\Big{)}$
@@
The average number of iterations required for convergence for each $N$ is shown in Table III. The largest mean number of iterations for an $M=16$, $N=128$ system occurs when $d_{r}=0.5$ and is 6.60. In this scenario, when the maximum number of iterations is fixed at 7, the resulting SNR averages to be 99.8% of the fully converged SNR. Hence, restricting the number of iterations to a fixed amount has little impact on performance.
TABLE III: Mean number of iterations of the ISD required for convergence
N Mean number of iterations
$\mathbf{d_{r}=0.1}$ $\mathbf{d_{r}=0.5}$
$N=4$ 2.40 2.35
$N=8$ 2.65 2.56
$N=16$ 2.96 2.96
$N=32$ 3.50 3.96
$N=64$ 4.12 5.04
$N=128$ 5.42 6.60
@@ V-B Mean Sum Rate of SD versus Non-Subsurface Methods
Many existing MU RIS phase selection methods, such as those in [5, 6, 7], focus on achieving near optimal performance. Necessarily, they have challenging computational complexity requirements. The lowest complexity MU method we are aware of is that first proposed in [4]. While other methods may offer better performance, the design in [4] is more likely to be considered for practical implementation. This method approximately minimises the total mean squared error (TMSE) of a MMSE receiver. All users are placed in one band, so each user has $K$ times the bandwidth (BW) when compared to users served by the SD. In [12], we compared the SD and TMSE method in terms of BW, complexity and fairness. The SD offers a reduction in complex multiplies by a factor of $N(K^{2}+K(N+2))$, requires no matrix operations, provides fairer results according to Jain’s fairness test, uses a much simpler receiver type (MF vs MMSE) and reduces CSI requirements by a factor of $K$. Hence, even when compared to the simplest RIS designs, the SD approach offers substantial complexity savings.
In this section, the mean sum rates for the SD and TMSE methods are compared with the mean sum rate for randomly selected phases. The inclusion of random phases creates a lower bound to compare both methods against. For the random approach, the phase shift of each element is generated from a uniform distribution between 0 and $2\pi$. It is assumed that all users are in one frequency band like the TMSE example. We investigate the impact of important system features, such as K-factor levels, correlation, system size and user locations.
@@ V-B1 K-Factors and Clustering
Figs. 4, 5 and 6 compare the mean sum rate for the SD with that of random phases and the TMSE design for varying levels of the UE-BS, RIS-BS and UE-RIS K-factor, respectively. $\eta$, as defined in (LABEL:eq:etad), (LABEL:eq:etarb) and (LABEL:eq:etaur), rather than K-factor itself is plotted so the proportion of the link that is LoS can be easily observed.
The array parameters for Figs. 4, 5 and 6 are $M\!=\!16$, $N\!=\!128$, $K\!=\!4$, $M_{x}\!=\!4$, $N_{x}\!=\!16$, $d_{b}\!=\!0.5$, and $d_{r}\!=\!0.1$. The two K-factors not being varied in each plot are set to 1 so that the LoS and NLoS component strengths are equal. The coloured lines designate the type of RIS design, while line types (given in black in the legend) designate the user locations.
@@
It can be seen in Figs. 4, 5 and 6 that despite the significant BW restriction on the SD, there are still operating scenarios where it achieves higher rates than the TMSE design for significantly lower complexity. It is also clear that a more dominant LoS link is beneficial for the SD, but negatively impacts the TMSE design. As the K-factor, and therefore $\eta$ (as defined in (LABEL:eq:etad), (LABEL:eq:etarb) and (LABEL:eq:etaur)) of a channel increases, the rank of that channel decreases until it reaches rank-1 when $\eta=1$. This limits the spatial multiplexing ability of the channel by reducing diversity. Insufficient diversity leads to the inability to separate MU channels. As all users are in separate bands for the SD, a lack of channel diversity is not an issue. The SNR improves as channels can be aligned so that more power is directed towards the user as the rank decreases. However, for the TMSE design where all users are located in one frequency band, the inability to separate channels as the K-factor increases leads to a significant drop in rate. A slight exception to this trend can be seen in Fig. 5 between $\eta_{\mathrm{rb}}=0$ and $\eta_{\mathrm{rb}}=0.3$. Note that the TMSE method assumes a LoS component in $\mathbf{H}_{\mathrm{rb}}$, so increasing the K-factor when $\eta_{\mathrm{rb}}$ is low helps slightly, as the channel becomes more similar to that which the framework was built upon. However, above this threshold, the reduction in diversity dominates and the overall rate reduces. The inability to separate channels also negatively impacts the random design, but this impact is lower as it is already operating at a low rate.
@@
Varying the K-factors for the UE-RIS and RIS-BS channels results in similar mean sum rate behaviour. Data through the RIS must travel through both channels, so changing the diversity of one places limits on the other, and thus on the whole RIS link. This is particularly prominent when the LoS link is dominant - the mean sum rate for the TMSE method drops dramatically, due to the direct link providing the only diversity. The same can be observed for the random method, but to lesser extent. In comparison, when the UE-BS channel becomes strongly LoS, the mean sum rate for the TMSE design is less impacted. As discussed at the start of Sec. V, UEs are dropped in locations that ensure a strong RIS path, to ensure the effects of the RIS are visible. Therefore, the rate remains high in this case due to the strong RIS link being unaffected.
@@
The impact of placing all 4 users in a cluster (layout A) is most noticeable in strongly LoS channels, where it improves the rate of the SD and lowers the rate of the TMSE method. Clustering users reduces channel diversity, and therefore negatively impacts the rate of the single frequency TMSE method. A key benefit of the SD is its robustness to clustered users, a realistic scenario in highly populated areas. As it does not rely on spatial multiplexing, clustered users improve its performance due to the similarity of the LoS paths, allowing those subsurfaces designed for other users to provide better signal enhancement. Interestingly, clustering also helps the random method. This is likely due to the fact that when the random phases happen to be beneficial for one user, they benefits all UEs in the cluster.

### v2

@@ IV-C Special Case 2: Mean SNR for LoS $\mathbf{H}_{\mathrm{rb}}$ with Correlated Rayleigh UE Channels
[…]
@@
Remark 8
As in Remark 7, alignment between $\mathbf{R}_{\mathrm{d}}^{(k)}$ and $\mathbf{a}_{\mathrm{b}}^{(k)}$ and increased spatial correlation tends to increase the SNR components in (33)-(35).
@@ V Numerical Results
It is assumed that the RIS is fixed at 40 m away from the BS, at an angle of $\frac{\pi}{4}$ rad. Users are then dropped in a straight corridor of width $2w$ between BS and RIS, at a maximum distance of $r$ from the RIS with an exclusion zone of 1 m around the RIS. This region ensures that the path from the UE through the RIS significantly contributes to the total channel, and the effects of the RIS can be clearly seen. User clusters are considered to investigate the effect of people naturally congregating. Three types of user grouping were used in simulations. As seen in Fig. 2, layout A places all $K$ UEs in one cluster, where they are all in the same area spaced 1 m apart and layout B creates 2 clusters of $\frac{K}{2}$ UEs, with users spaced 1 m apart. Cluster locations are drawn from a uniform distribution. Layout C places each of the $K$ UEs randomly within the corridor, again using a uniform distribution. To ensure that results are general and not tied to a specific UE location, $10^{4}$ random user drops are generated per layout, each with $10^{6}$ replicates.
@@
The steering vectors $\mathbf{a}_{\mathrm{d}}$, $\mathbf{a}_{\mathrm{b}}$, $\mathbf{a}_{\mathrm{r}}$ and $\mathbf{a}_{\mathrm{ur}}$ correspond to the vertical uniform linear array (VURA) model and are given in [28]. Elements are arranged in a VURA at intervals of $d_{b}$ and $d_{r}$ wavelengths at the BS and RIS, respectively. $M_{x}$ and $N_{x}$ are the number of BS and RIS elements per row and $M_{z}$ and $N_{z}$ are the number of BS and RIS elements per column, such that $M=M_{x}M_{z}$ and $N=N_{x}N_{z}$. $\theta_{\mathrm{D}}$ and $\phi_{\mathrm{D}}$ are the elevation and azimuth angles-of-departure (AoDs) at the RIS and $\theta_{\mathrm{A}}$ and $\phi_{\mathrm{A}}$ are the corresponding elevation and azimuth angles-of-arrival (AoAs) at the BS. We assume the RIS is on a $\frac{\pi}{4}$ rad angle with respect to the BS, so $\phi_{\mathrm{D}}=\frac{5\pi}{4}$ rad and $\phi_{\mathrm{A}}=\frac{\pi}{4}$ rad. We also assume both are at the same height, so $\theta_{\mathrm{D}}=\theta_{\mathrm{A}}=\frac{\pi}{2}$ rad.
Without loss of generality we assume $\sigma^{2}=1$ and $E_{s}$ is selected in Sec V-A so that the mean output SNR for one UE is 5 dB in a system using the SD where $N=128$, $M=16$, $d_{r}=0.1$, $d_{b}=0.5$ and $\kappa_{d}=\kappa_{rb}=\kappa_{ur}=1$. In Sec. V-B, we select $E_{s}$ so that the mean output SNR for one UE is 5 dB in a system using the SD where $N=128$, $M=16$, $d_{r}=0.1$, $d_{b}=0.5$, $\kappa_{\mathrm{ur}}=\kappa_{\mathrm{d}}=1$ and the RIS-BS channel is purely LoS, an important special case. These parameter values and definitions do not change throughout the results, unless specified otherwise.
@@
Fig. 3 verifies the analytical results in Sec. IV-A by showing the equivalence of $\mathbb{E}[\mathrm{SNR}]$ and the simulated mean SNR for the SD, and then compares this with the simulated mean SNR for the ISD. Two versions of the ISD were implemented. The first iterates through all subsurfaces once, as shown in Sec. III-C. The second iterates through all subsurfaces and calculates the SNR of the system. It then repeats this process multiple times until the SNR of the system has converged, i.e. it increases by less than a specified tolerance from one iteration to the next. For these simulations, the tolerance was specified to be $10^{-4}$. This method is referred to as the CISD. The array parameters for Fig. 3 are $M=16$, $M_{x}=4$, $d_{b}=0.5$ and all K-factors are 1.
@@
Figure 3: Comparison of mean SNR for a randomly selected UE for different subsurface phase selection methods while varying the number of RIS elements.
@@
It can be seen that the simulated mean SNR for the method outlined in Sec. III-A is virtually indistinguishable from the $\mathbb{E}[\mathrm{SNR}]$ derived in Sec. IV-A, verifying that result. As expected, the mean SNR for all 3 methods increases with $N$, however the rate of improvement decreases as $N$ increases. This is expected, as it is well known that the RIS SNR can grow quadratically with $N$ [18] so that the dB relationship is logarithmic. The quadratic behavior can be seen in (23), (31), and (34) where the $\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}]$ term contains the sum of $N_{k}(N_{k}-1)$ positive terms, giving a contribution of $\mathcal{O}(N_{k}^{2})$. The ISD and converged ISD both provide more benefit at higher $N$. At high $N$, more improvement can be made by iterating, as there are more elements to vary and hence more opportunities to improve the performance of the system. This improvement is more pronounced when the correlation between RIS elements is lower. Figure 3 shows that the gap between the mean SNR of the SD and CISD is larger for RIS element spacings of $d_{r}=0.5$ than for $d_{r}=0.1$. This may be due to the fact that there is more variability in the previously set subsurfaces creating the fixed channels for the iterative design. Hence, reduced correlation leads to greater variation in the fixed channels, which the iterative design can exploit. Additionally, when $N$ is kept constant, the average SNR per UE decreases as the number of UEs increases. This reduction arises because the fixed set of RIS elements must be divided between more UEs, resulting in fewer elements being allocated to each UE.
A key design choice for the ISD is the order in which subsurfaces are set. The standard ISD in Sec. III-C sets the subsurface of weaker UEs first. This allows the strongest UEs to benefit from additional channel information, and leads to higher average performance. However, this does not lead to fair performance across all UEs. Table I compares the ISD to two other versions. The first is a reverse version, where strongest UEs are set first, allowing the weakest UEs to benefit from extra information. The second is a random version, where the order in which the UEs are set is random. This version acts as a benchmark and allows insights into the respective benefits provided by the ISD and reverse ISD. The mean SNR of the strongest UE (UE 1) to the weakest UE (UE 4) for all three methods is shown for $N=128$ and $d_{r}=0.5$.
TABLE I: Mean SNR ($\mathrm{dB}$) for each UE using the ISD and the reverse ISD.
UE 1 UE 2 UE 3 UE 4
ISD 15.12 7.87 3.35 0.30
Reverse ISD 13.38 7.25 3.44 0.93
Random ISD 14.42 7.51 3.35 0.63
The reverse ISD offers fairer performance at the expense of a rate drop for the stronger UEs and a lower average rate. The random ISD provides a balance between the two, with less of a rate drop than the reverse ISD, but also lower fairness. Thus, the choice of subsurface ordering should depend on whether fairness or average rate is the priority.
@@
Multiplies KM+NKM\!+\!N $NK\!+\!n\!\Big(\!\frac{(K\!-\!1)!MN}{K}$ $\times\!\!\left(\!\frac{N}{K}\!+\!1\!\right)\!+\!KM\!+\!N\!\Big)$ $NK\!+\!n\!\Big(\!\frac{(K\!-\!1)!MN}{K}$ $\times\!\!\left(\!\frac{N}{K}\!+\!1\!\right)\!+\!KM\!(2\!+\!N(N\!+\!1)\!)\!+\!N\Big)$
@@
The average number of iterations required for convergence for each $N$ is shown in Fig. 4. The largest mean number of iterations for an $M=16$, $N=128$ system occurs when $d_{r}=0.5$ and is 6.60. In this scenario, when the maximum number of iterations is fixed at 7, the resulting SNR averages to be 99.8% of the fully converged SNR. Hence, restricting the number of iterations to a fixed amount has little impact on performance.
Figure 4: Mean number of iterations of the ISD required for convergence.
@@ V-B Mean Sum Rate of SD versus Non-Subsurface Methods
Many existing MU RIS phase selection methods, such as those in [4, 8, 7, 10], focus on achieving near optimal performance. Necessarily, they have challenging computational complexity requirements. The lowest complexity MU method we are aware of is that proposed in [9]. While other methods may offer better performance, the design in [9] is more likely to be considered for practical implementation. This method approximately minimizes the total mean-squared error (TMSE) of a minimum mean-squared error (MMSE) receiver. All users are placed in one band, so each user has $K$ times the bandwidth (BW) when compared to users served by the SD. This discrepancy in available BW per UE would exist between the SD and any MU approach where all of the UEs are in a single frequency band. We believe it is important to compare our method to this type of scenario, as this is the most common type of MU RIS design in the literature. Also, considering a key benefit of the SD is a significant complexity reduction, a low-complexity MU approach was considered to be a fair comparison. To show the low complexity nature of the TMSE method and the significant complexity reduction afforded by the SD, the number and complexity of matrix operations required for both methods is compared to the low-complexity near-optimal approach in [10] in Table III. The approach in [10] jointly optimizes the RIS phase shifts, precoder and combiner for a single UE system to minimize the MSE. For a fair comparison with the SD and TMSE approaches, the number of operations for the single UE case are scaled by $K$, representing an OFDM-type system with no IUI.
TABLE III: Complexity comparison of the SD, the TMSE method in [9] and the MSE method in [10].
SD TMSE [9] MSE [10]
Complex Multiplicative Operations 3N+K(M+1)3N\!+\!K(M\!+\!1) $(N\!+\!M)(K\!+\!K^{2})\!+\!K(3\!+\!3K+\!2K^{2})$ nK(2N(M+1)+M+3)nK\!(2N(M\!+\!1)\!+\!M\!+\!3)
Matrix Functions None 2 $K\!\times\!K$ inverses
1 $K\!\times\!K$ eigendecomposition 1 $M\times 1$ eigendecomposition
Assuming that $N$ is the largest variable, the SD offers a multiplicative reduction of a factor of $K+K^{2}$ over the TMSE method, and of $nKM$ over the MSE approach. For a system where $N=128$, $M=32$, $K=4$ and $n=40$ (as suggested in [10]), the SD, TMSE method and MSE method require 516, 3388 and 1356960 multiplicative operations, respectively, highlighting the significant computational savings provided by the SD. Additionally, the SD does not require matrix functions, unlike the others. Results in [16] also showed that the SD provides fairer results according to Jain’s fairness test, uses a much simpler receiver type (MF vs MMSE) and reduces CSI requirements by a factor of $K$ when compared to the TMSE approach. Hence, the SD approach offers substantial complexity savings even compared to the simplest RIS designs.
In this section, the mean sum rates for the SD and TMSE methods are compared with the mean sum rate for randomly selected phases. The inclusion of random phases creates a lower bound to compare both methods against. For the random approach, the phase shift of each element is generated from a uniform distribution between 0 and $2\pi$. It is assumed that all users are in one frequency band like the TMSE example. We investigate the impact of important system features, such as K-factor levels, correlation, system size and UE locations.
@@ V-B1 K-Factors and Clustering
Figs. 5, 6 and 7 compare the mean sum rate for the SD with that of random phases and the TMSE design for varying levels of the UE-BS, RIS-BS and UE-RIS K-factor, respectively. $\eta$, as defined in (LABEL:eq:etad), (LABEL:eq:etarb) and (LABEL:eq:etaur), rather than K-factor itself is plotted so the proportion of the link that is LoS can be easily observed.
The array parameters for Figs. 5, 6 and 7 are M=16M\!=\!16, N=128N\!=\!128, K=4K\!=\!4, $M_{x}\!=\!4$, $N_{x}\!=\!16$, $d_{b}\!=\!0.5$, and $d_{r}\!=\!0.1$. The two fixed K-factors in each plot are set to 1 so that the LoS and NLoS component strengths are equal. Note that RLU represents randomly located UEs and 1 UC and 2 UCs represent one and two clusters of UEs, respectively.
@@
Remarkably, Figs. 5, 6 and 7 show that despite the significant BW restriction and lower complexity, the SD actually outperforms the TMSE when the K-factor of one of the RIS links is high. A LoS or near-LoS RIS link is a key operating scenario that would be deliberately selected in practice. RIS are often motivated as a method of overcoming propagation challenges in the direct path, and therefore, it is expected that their placement could be selected to ensure a strong LoS path to the BS and/or UEs. In such configurations, a high K-factor is not only realistic but desirable. Thus, the SD excels in a scenario of significant practical relevance and is well positioned for use in deployed RIS systems. While the TMSE method yields higher sum rates at lower K-factors, its complexity scales much more rapidly with system size than that of the SD. Therefore, even for lower K-factor values, the SD has benefits for larger systems.
@@
A higher K-factor is beneficial for the SD but negatively impacts the TMSE design. As the K-factor, and therefore $\eta$ (as defined in (LABEL:eq:etad), (LABEL:eq:etarb) and (LABEL:eq:etaur)) of a channel increases, the rank of that channel decreases until it reaches rank-1 when $\eta=1$. This limits the spatial multiplexing ability of the channel by reducing diversity. Insufficient diversity leads to the inability to separate MU channels. As all users are in separate bands for the SD, a lack of channel diversity is not an issue. The SNR improves as channels can be aligned so that more power is directed towards the user as the rank decreases. However, for the TMSE design where all users are located in one frequency band, the inability to separate channels as the K-factor increases leads to a significant drop in rate. A slight exception to this trend can be seen in Fig. 6 between $\eta_{\mathrm{rb}}=0$ and $\eta_{\mathrm{rb}}=0.3$. Note that the TMSE method assumes a LoS component in $\mathbf{H}_{\mathrm{rb}}$, so increasing the K-factor when $\eta_{\mathrm{rb}}$ is low helps slightly, as the channel becomes more similar to that which the framework was built upon. However, above this threshold, the reduction in diversity dominates and the overall rate reduces. The inability to separate channels also negatively impacts the random design, but this impact is lower as it is already operating at a low rate.
@@
Varying the K-factors for the UE-RIS and RIS-BS channels results in similar mean sum rate behaviour. Data through the RIS must travel through both channels, so changing the diversity of one places limits on the other, and thus on the whole RIS link. This is more prominent when the LoS link is dominant - the mean sum rate for the TMSE method drops dramatically, due to the direct link providing the only diversity. The same can be observed for the random method, but to lesser extent. In comparison, when the UE-BS channel becomes strongly LoS, the mean sum rate for the TMSE design is less impacted. As discussed at the start of Sec. V, UEs are dropped in locations that ensure a strong RIS path, to ensure the effects of the RIS are visible. Therefore, the rate remains high in this case due to the strong RIS link being unaffected.
@@
The impact of placing all four users in a cluster (layout A) is most noticeable in strongly LoS channels, where it improves the rate of the SD and lowers the rate of the TMSE method. Clusters of densely located UEs are expected in realistic scenarios of highly populated areas, such as stadiums or transit hubs. Clustering users reduces channel diversity and negatively impacts the rate of the single frequency TMSE method. A key benefit of the SD is its robustness to clustered users. As it does not rely on spatial multiplexing, clustered users improve its performance due to the similarity of the LoS paths, allowing those subsurfaces designed for other users to provide better signal enhancement. Interestingly, clustering also helps the random method. This is likely due to the fact that when the random phases happen to be beneficial for one user, they benefits all UEs in the cluster.